OpenStudy (anonymous):

If h(x) = −5x − 10, find h(−2)

OpenStudy (anonymous):

h(-2) = -5(-2) - 10 = 10 - 10 = 0. 0 is the answer.

OpenStudy (anonymous):

Substitute -2 in place of x's. So h(-2) = -5(-2) -10. h(-2) = --10 -10. double negatives is +. so 10-10 = 0.